# Spatial Phonons: A Phenomenological Viscous Dark Energy Model for DESI

### v2

@@ 4 Cosmological Implications and Comparison with DESI
For this, first recall that the effective dark energy equation of state in the one–channel phonon model is
$w_{\rm eff}(H)=-1+\varepsilon-3\kappa\,f\left(x_{\rm ph}(H)\right),\quad f(x)\equiv\frac{x}{1+x^{2}},\quad x_{\rm ph}(H)\equiv H\,\tau_{\rm ph}(H)$ (4.1)
For appropriate choices of $\varepsilon$, $\kappa$, $H_{\star}$ and the parameters that control $\tau_{\rm ph}(H)$, equation (4.1) can reproduce the qualitative behaviour favoured by DESI–based reconstructions of the expansion history. For example, choosing $\varepsilon$ and $\kappa$ slightly below unity and taking the mass–gap scale $H_{\star}$ such that $x_{\rm ph}(H)$ passes through unity at a redshift $z_{\star}$ of order unity, one can obtain a $w_{\rm eff}(z)$ that lies close to Chevallier–Polarski–Linder (CPL) fits extracted from DESI BAO in combination with Planck CMB and Type Ia supernovae data [20, 24, 52, 50, 58, 47]. In such fits the effective equation of state is mildly above w=1w=-1 at $z\simeq 0$, crosses the phantom divide at moderate redshift and dips to more negative values around $z\sim 1$, before gradually relaxing again as the relative dark energy density falls at higher redshift.
For concreteness, we now make contact with DESI–based CPL fits in a practical way and fix the parameters $\varepsilon$, $\kappa$ and $H_{\star}$. In Sec. 3.3 we found that the dimensionless relaxation coefficient takes the form,
@@
where $H_{\star}$ is the Hubble parameter that sets the phonon mass gap. We also assume that the Gibbons–Hawking temperature $T_{H}$ at the epoch of dark energy domination is already below the critical temperature $T_{H_{\star}}$.
[…]
$H_{\Lambda}^{2}(z)=H_{0}^{2}\left[\Omega_{m0}(1+z)^{3}+1-\Omega_{m0}\right]$ (4.3)
with $\Omega_{m0}\simeq 0.3$ the present non–relativistic matter fraction. Using this approximate background we match,
$w_{\rm eff}(H)=-1+\varepsilon-3\kappa\,\frac{x_{\rm ph}(H)}{1+x_{\rm ph}(H)^{2}}$ (4.4)
to a representative CPL form that encodes the DESI preference for mild evolution in the dark energy sector. Current combinations of Planck CMB, CMB lensing, DESI BAO and Type Ia supernova data provide tight constraints on the Chevallier–Polarski–Linder (CPL) parameters $(w_{0},w_{a})$. For example, Chudaykin and Kunz [51] find,
$w_{0}=-0.856\pm 0.062,\qquad w_{a}=-0.53^{+0.28}_{-0.26}\quad(68\%~\text{C.L.})$ (4.5)
for a combined analysis of Planck, DESI DR1 BAO and supernova data in a $w_{0}$$w_{a}$CDM model.
The CPL form [20, 24] reads,
$w_{\rm CPL}(z)=w_{0}+w_{a}\frac{z}{1+z}$ (4.6)
which gives the following CPL equations of state at the redshifts z=0, 0.5, 1.0,1.1, 1.2, 1.3, 1.4z=0,\ 0.5,\ 1.0,\\ 1.1,\ 1.2,\ 1.3,\ 1.4 and $1.5$ for the Chudaykin–Kunz pair (4.5),
$z$ $w_{\rm CPL}(z)$
0.0 0.856-0.856
0.5 1.033-1.033
1.0 1.121-1.121
1.1 1.134-1.134
1.2 1.145-1.145
1.3 1.156-1.156
1.4 1.165-1.165
1.5 1.174-1.174
A simple least–squares fit that matches $w_{\rm eff}$ in equation (4.4) to $w_{\rm CPL}(z)$ at these redshifts yields the parameter set,
$\varepsilon\simeq 0.335,$
$\kappa\simeq 0.3349,$
$\frac{H_{\star}}{H_{0}}\simeq 2.10$ (4.7)
for which the phonon sound speed is almost luminal,
$c_{s}^{2}=\frac{\kappa}{\varepsilon}\simeq 0.9997$ (4.8)
With this choice the function $x_{\rm ph}(H)$ passes through unity at $H(z_{\star})\simeq H_{\star}$, which corresponds to a redshift $z_{\star}\simeq 1.31$ in the approximate background (4.3), that is, in the middle of the DESI galaxy–BAO redshift range where the constraints on the expansion history are strongest.
Using the above choice of parameters we obtain the effective equation of state values as,
$z$ $w_{\rm eff}(z)$
0.0 0.820-0.820
0.5 0.971-0.971
1.0 1.135-1.135
1.1 1.153-1.153
1.2 1.163-1.163
1.3 1.167-1.167
1.4 1.165-1.165
1.5 1.158-1.158
These values show that the phonon model can reproduce both the mild deviation from w=1w=-1 at $z=0$ and the transient phantom dip around $z\sim 1$ suggested by CPL fits to the DESI–based datasets, while keeping $\kappa$ and $\varepsilon$ sub–unity and maintaining a nearly luminal phonon sound speed. The difference between $w_{\rm eff}$ and $w_{\rm CPL}$ is at most a few percent as summarized in table 2. In addition to this we see that the model again starts returning to a non phantom behavior for larger redshifts ($z\gtrsim 1.5$) whereas the extrapolation of the DESI priors result in stronger phantoms in the past. It is therefore useful to understand what the phonon model predicts for the very early and very late universe compared to DESI-CPL extrapolations.
The key quantity is the dimensionless relaxation parameter $x_{\rm ph}(H)=H\tau_{\rm ph}(H)$ in (4.1). For $H\gg H_{\star}$ one has
$x_{\rm ph}(H)=\frac{H_{\star}}{H}\,\exp\left(\frac{H_{\star}}{H}-1\right)\;\approx\;\frac{H_{\star}}{e\,H}\quad\text{for}\quad H\gg H_{\star}$ (4.9)
so $x_{\rm ph}(H)\to 0$ in the very early universe. For $H\ll H_{\star}$ the factor $H_{\star}/H$ becomes very large and the exponential dominates, so $x_{\rm ph}(H)\to\infty$ in the infinite future. In both limits the Maxwell kernel in (4.1) satisfies,
$\frac{x_{\rm ph}(H)}{1+x_{\rm ph}(H)^{2}}\longrightarrow 0\quad\text{as}\quad H\to\infty\quad\text{or}\quad H\to 0$ (4.10)
and the effective equation of state approaches the simple asymptotic value
$\lim_{H\to\infty}w_{\rm eff}(H)=\lim_{H\to 0}w_{\rm eff}(H)=-1+\varepsilon$ (4.11)
In our least–squares fit to the DESI–motivated CPL targets the phonon parameters are driven close to,
$\kappa\ \simeq\ \varepsilon\ \simeq\ \frac{1}{3}\ \simeq\ 0.33$ (4.12)
which implies a phonon sound speed $c_{s}^{2}=\kappa/\varepsilon$ very close to unity and an asymptotic dark energy equation of state,
$w_{\rm eff}\ \longrightarrow\ -1+\varepsilon\ \simeq\ -\frac{2}{3}\ \simeq\ -0.66$ (4.13)
@@
both in the early universe and in the infinite future. The bulk viscous term in (4.1) is therefore a genuinely transient effect that operates only while $x_{\rm ph}(H)$ is of order unity, that is, while the Hubble rate is comparable to the mass gap scale $H_{\star}$. In this intermediate regime the effective equation of state dips below w=1w=-1 and tracks the DESI–driven CPL reconstruction, while at very early and very late times the model relaxes back to a regular quintessence–like value $w_{\rm eff}\simeq-2/3$.
For comparison, a naive extrapolation of the CPL fit with the Chudaykin–Kunz parameters (4.5) implies $w(z)\to w_{0}+w_{a}\simeq-1.4$ in the infinite past and $w(z)\to+\infty$ as $z\to-1$, even though the dark energy density itself tends to zero in both limits [51]. The phonon model instead has finite and equal limits in the distant past and future and the phantom phase is localized near $z_{\star}\sim 1.3$, which is the regime directly probed by DESI.
It is worth noting that the DESI motivated CPL fit selects a parameter set that is quite natural from the point of view of the viscoelastic model. The values $\kappa\simeq\varepsilon\simeq 1/3$ are less than unity, as expected, and both represent a moderate fraction of the brane tension $T_{s}$. At the same time the associated phonon sound speed, $c_{s}^{2}=\kappa/\varepsilon\simeq 1$, lies very close to the causal limit without overshooting it. This is a plausible outcome for longitudinal modes in a very stiff medium and it shows that the DESI driven fit does not require an exotic or finely tuned region of parameter space. Instead fitting DESI using our viscous dark energy model naturally singles out a regime in which the brane behaves like a strongly elastic but still well behaved effective medium where the phonon enthalpy and bulk modulus are each of order one third of the geometric tension $T_{s}$ and the phonons are naturally causal. As such, given below are tables summarizing the findings and a comparison plot.
Quantity Symbol Value
Phonon enthalpy fraction $\varepsilon$ $0.335$
Stiffness parameter $\kappa$ $0.33489$
Mass gap scale $H_{\star}/H_{0}$ $2.10$
Sound speed squared $c_{s}^{2}=\kappa/\varepsilon$ $0.9997$
Table 1: Example parameter choice for the one channel phonon model which reproduces a CPL like dark energy equation of state in the DESI redshift range. Note that the ‘mass gap’ of $2.10$ implies that the phonons are approximately $2.10$ times heavier than the present day Hubble scale.
Redshift $z$ $w_{\rm CPL}(z)$ $w_{\rm eff}(z)$ $|w_{\rm eff}-w_{\rm CPL}|$
0 0.856-0.856 0.820-0.820 $0.036$
$0.5$ 1.033-1.033 0.971-0.971 $0.062$
$1.0$ 1.121-1.121 1.135-1.135 $0.014$
$1.1$ 1.134-1.134 1.153-1.153 $0.019$
$1.2$ 1.145-1.145 1.163-1.163 $0.018$
$1.3$ 1.156-1.156 1.167-1.167 $0.011$
$1.4$ 1.165-1.165 1.165-1.165 $0.000$
$1.5$ 1.174-1.174 1.158-1.158 $0.016$
Table 2: CPL target equation of state and the prediction of the one channel phonon model with parameters in Table 1. The difference is at the few percent level over the entire DESI sensitive redshift range.
Figure 1: Comparison between the Chevallier–Polarski–Linder equation of state $w_{\rm CPL}(z)$ (solid line) with parameters $(w_{0},w_{a})=(-0.856,-0.53)$ and the effective equation of state $w_{\rm eff}(z)$ (dashed line) of the viscoelastic spatial phonon model for $\varepsilon=0.335$, $\kappa=0.33489$ and $H_{\star}/H_{0}=2.10$. The horizontal dotted line shows the cosmological constant value w=1w=-1. Over the DESI redshift range $0\leq z\lesssim 1.6$, the viscous dark energy model closely tracks the CPL fit.
We now close by commenting on the mass scale of the phonons.
In the simplified description of Section 3 we modeled the relaxation time as a function of the Hubble rate and introduced a characteristic scale $H_{\star}$ at which the viscous response is strongest. We identified this scale with the phonon mass through,

### v3

@@ 4.3 Model Prediction $w_{\rm eff}(z)$ and mapping to an Effective CPL Pair
Recall that the phonon model predicts an effective dark energy equation of state as a function of the Hubble rate,
$w_{\rm eff}(H)\;=\;-1\,+\,\varepsilon\,-\,3\,\kappa\,\frac{x_{\rm ph}(H)}{1+x_{\rm ph}^{2}(H)}\,,\qquad x_{\rm ph}(H)\;\equiv\;H\,\tau_{\rm ph}(H)$ (4.3)
The ancillary script uses the same relaxation model as the main text, written in dimensionless form with $h\equiv H/H_{0}$ and $h_{\star}\equiv H_{\star}/H_{0}$ and therefore,
$x_{\rm ph}(h)\;=\;\frac{h_{\star}}{h}\exp\left(\frac{h_{\star}}{h}-1\right)$ (4.4)
[…]
$h^{2}_{\Lambda}(z)\;\equiv\;\left(\frac{H_{\Lambda}(z)}{H_{0}}\right)^{2}\;=\;\Omega_{m0}(1+z)^{3}+(1-\Omega_{m0})\,,\qquad\Omega_{m0}=0.3$ (4.5)
We then evaluate $w_{\rm eff}(z)$ by composing equations (4.3) to (4.5).
To compare the model to the CPL plane we fit the model prediction to the CPL form (Chevallier & Polarski (2001); Linder (2003)),
$w_{\rm CPL}(z)\;=\;w_{0}\,+\,w_{a}\,\frac{z}{1+z}$ (4.6)
Define $y(z)\equiv z/(1+z)$ so that $w_{\rm CPL}(z)=w_{0}+w_{a}\,y(z)$. Over a dense grid $0\leq z\leq 1.6$ with $N=600$ points, the script performs an ordinary least squares regression of $w_{\rm eff}(z)$ onto the basis $\{1,y(z)\}$ and returns the effective CPL pair,
$\theta_{\rm mod}(\varepsilon,\kappa,h_{\star})\;\equiv\;\begin{pmatrix}w^{\rm mod}_{0}\\ w^{\rm mod}_{a}\end{pmatrix}$ (4.7)
4.4 Compressed Gaussian Likelihood and $\chi^{2}$ Statistic
We approximate the posterior in the $(w_{0},w_{a})$ plane by a bivariate Gaussian described by equation (4.2). For a model prediction $\theta_{\rm mod}$ we define,
$\chi^{2}(\varepsilon,\kappa,h_{\star})\;=\;\Delta\theta^{\,T}\,\mathbf{C}^{-1}\,\Delta\theta\,,\qquad\Delta\theta\;=\;\theta_{\rm mod}-\hat{\theta}$ (4.8)
This quantity is the squared Mahalanobis distance between the model implied CPL pair and the observational mean. In our analysis we scan over internal model parameters to locate points whose CPL image lies close to the DESI mean, rather than performing a full likelihood fit to the underlying BAO, supernova, and CMB data. Therefore a very small value $\chi^{2}\ll 1$ should be read as a consistency and viability statement at the level of the compressed summary, not as a conventional goodness of fit with an associated $p$ value.
The parameter scan in Statistical Fit.py imposes the basic physical prior,
$0\;<\;\kappa\;\leq\;\varepsilon$ (4.9)
The condition $\varepsilon\geq\kappa$ together with $\varepsilon<1$ and $\kappa<1$ prevents tachyonic behavior in the effective phonon description and is reasonable because $\varepsilon$ and $\kappa$ are defined as fractional response coefficients, so values above unity would correspond to an unphysical over response in a minimal model. A definitive statistical assessment requires a self consistent solution for $H(z)$ within the model and a direct fit to the full likelihoods, which we leave to future work. The scan searches within the box,
$0.05\leq\varepsilon\leq 0.95\,,\qquad 0.05\leq\kappa\leq 0.95\,,\qquad 0.5\leq h_{\star}\leq 4.0$ (4.10)
A coarse grid scan is followed by local refinements around the best point. This design keeps the analysis fast and fully reproducible.
4.5 Numerical Results
For the DESI DR1 compressed summary in equation(4.2), the ancillary script finds the best fit,
$(\varepsilon,\kappa,h_{\star})_{\rm best}\;=\;(0.596835,\;0.574051,\;1.794118)$ (4.11)
which maps to,
$(w^{\rm mod}_{0},w^{\rm mod}_{a})_{\rm best}\;=\;(-0.829119,\;-0.744835)\,,\qquad\chi^{2}_{\rm best}\;=\;1.02088\times 10^{-3}$ (4.12)
It is also useful to report the implied phonon sound speed diagnostic,
$c_{s}^{2}\;\equiv\;\frac{\kappa}{\varepsilon}$ (4.13)
which for the best fit gives,
$c_{s}^{2}\;\simeq\;0.961825\,,\qquad\frac{\varepsilon-\kappa}{\varepsilon}\;\simeq\;0.038175$ (4.14)
No additional near luminal band is imposed in the scan beyond $\kappa\leq\varepsilon$. The closeness of $c_{s}^{2}$ to unity is therefore an output of the fit rather than a built in restriction.
Case $w_{0}$ $w_{a}$ $\Delta w_{0}\equiv w_{0}-\hat{w}_{0}$ $\Delta w_{a}\equiv w_{a}-\hat{w}_{a}$
DESI compressed mean $\hat{\theta}$ 0.828209-0.828209 0.744750-0.744750 0 0
Best fit 0.829119-0.829119 0.744835-0.744835 $9.10\times 10^{-4}$ $8.45\times 10^{-5}$
Table 1: Comparison between the compressed CPL mean $(\hat{w}_{0},\hat{w}_{a})$ and the model implied CPL pair obtained from the regression procedure in Sec. 4.3.
Case $\varepsilon$ $\kappa$ $h_{\star}$ $c_{s}^{2}=\kappa/\varepsilon$ $(\varepsilon-\kappa)/\varepsilon$ $\chi^{2}$
Best fit $0.596835$ $0.574051$ $1.794118$ $0.961825$ $0.038175$ $1.02088\times 10^{-3}$
Table 2: Summary of the unrestricted best fit in the compressed $(w_{0},w_{a})$ likelihood. The scan imposes only $0<\kappa\leq\varepsilon$.
4.6 Pointwise Comparison of CPL Reconstructions and the model implied CPL Pair
In order to visualize how close the model implied CPL pair is to the compressed DESI mean in a more direct way, it is useful to compare the corresponding CPL equations of state at representative redshifts in the DESI leverage range. Given a CPL pair $(w_{0},w_{a})$ we define,
$\omega(z)\;=\;w_{0}\,+\,w_{a}\,\frac{z}{1+z}$ (4.15)
We evaluate equation (4.15) for the DESI compressed mean $(\hat{w}_{0},\hat{w}_{a})$ from equation (4.2) and for the model implied CPL pair $(w_{0}^{\rm mod},w_{a}^{\rm mod})$ from equation (4.12). Table 3 shows $\omega(z)$ and $\omega_{\rm eff}(z)$ at a set of representative redshifts $z=0.5$ and $z=1.0$ and in the interval $1.1\leq z\leq 1.5$ in steps of $0.1$. The final column reports the difference $\Delta\omega(z)\equiv\omega_{\rm eff}(z)-\omega(z)$. The near equality of the two columns reflects the fact that the best fit lies extremely close to the compressed mean in the $(w_{0},w_{a})$ plane.
$z$ $\omega(z)$ from $(\hat{w}_{0},\hat{w}_{a})$ $\omega_{\rm eff}(z)$ from $(w_{0}^{\rm mod},w_{a}^{\rm mod})$ $\Delta\omega(z)$
$0.5$ 1.076509-1.076509 1.077398-1.077398 $8.89\times 10^{-4}$
$1.0$ 1.200584-1.200584 1.201537-1.201537 $9.53\times 10^{-4}$
$1.1$ 1.218985-1.218985 1.219947-1.219947 $9.62\times 10^{-4}$
$1.2$ 1.234891-1.234891 1.235855-1.235855 $9.64\times 10^{-4}$
$1.3$ 1.248762-1.248762 1.249726-1.249726 $9.64\times 10^{-4}$
$1.4$ 1.260957-1.260957 1.261919-1.261919 $9.62\times 10^{-4}$
$1.5$ 1.271761-1.271761 1.272721-1.272721 $9.60\times 10^{-4}$
Table 3: Pointwise comparison of CPL equations of state. The second column uses the DESI compressed mean $(\hat{w}_{0},\hat{w}_{a})$, while the third column uses the model implied CPL pair $(w_{0}^{\rm mod},w_{a}^{\rm mod})$. The final column shows $\Delta\omega(z)\equiv\omega_{\rm eff}(z)-\omega(z)$.
Figure 1 provides the corresponding continuous comparison over $0\leq z\leq 1.5$. The two curves are visually indistinguishable on the plot, which is consistent with the small values of $\Delta\omega(z)$ in Table 3 and with the very small best fit $\chi^{2}$ reported in equation (4.12).
Figure 1: Overlay of the CPL equation of state $\omega(z)$ reconstructed from the DESI compressed mean $(\hat{w}_{0},\hat{w}_{a})$ and the model implied CPL equation of state $\omega_{\rm eff}(z)$ computed from $(w_{0}^{\rm mod},w_{a}^{\rm mod})$.
4.7 Interpretation and Limitations of the Compressed Approach
The extremely small value of $\chi^{2}_{\rm best}$ indicates that there exists a physically allowed region of $(\varepsilon,\kappa,h_{\star})$ for which the model implied CPL pair is essentially coincident with the mean of the compressed posterior in the $(w_{0},w_{a})$ plane. This should be interpreted as a consistency check that the phonon model can reproduce the central region of the published $w_{0}w_{a}$ constraints for the chosen dataset combination. It is worth emphasizing that the best fit lies in a physically natural regime for the phonon interpretation. The scan is allowed to explore a broad box of parameters, subject only to the basic causality prior $0<\kappa\leq\varepsilon$, yet it selects a point with $\kappa<\varepsilon<1$ and a small fractional splitting $\dfrac{\varepsilon-\kappa}{\varepsilon}\;\simeq\;0.038$, which corresponds to a near luminal phonon propagation speed $c_{s}^{2}=\kappa/\varepsilon\simeq 0.962$. This outcome is particularly appealing because near luminal propagation is precisely what one expects for very soft infrared oscillations of a stiff brane like medium. In our setup the characteristic scale $H_{\star}$ is of order the Hubble rate at redshift of order unity, so the associated phonon mass gap $m_{\phi}=H_{\star}/(2\pi)$ is extremely small in absolute units and only a fraction of $H_{0}$ (see below). The resulting Compton wavelength is therefore of order the Hubble horizon, meaning that the excitations relevant for the transient viscous response are horizon scale collective modes rather than microscopic degrees of freedom as mentioned earlier. For such long wavelength modes, relativistic propagation is the generic expectation unless the medium contains significant additional structure that enforces a much lower sound speed. In this sense the compressed likelihood fit does not merely accommodate near luminal phonons, it points to them. The parameter scan naturally identifies a region in which the brane is strongly elastic but still well behaved, with $\varepsilon$ and $\kappa$ both below unity and separated by only a few percent, so that the phonons remain causal while being very close to the luminal limit.
[…]
4.8 Asymptotic Equation of State of Dark Energy
Another important point concerns the asymptotic behaviour of the effective dark energy sector in the infinite past and in the infinite future in the context of the phonon model. A useful diagnostic of the far past and far future limits is provided by the dimensionless relaxation variable,
@@ 4.8 Asymptotic Equation of State of Dark Energy
In the asymptotic past, $H\gg H_{\star}$, we have $H_{\star}/H\to 0$ and therefore $x_{\rm ph}(H)\to 0$. In the asymptotic future, $H\ll H_{\star}$ so $x_{\rm ph}(H)\to 0$. In both limits the Maxwell kernel appearing in equation (4.3) satisfies,
$\frac{x_{\rm ph}(H)}{1+x_{\rm ph}^{2}(H)}\;\longrightarrow\;0\qquad\text{as}\qquad H\to\infty\quad\text{or}\quad H\to 0$ (4.17)
which implies that the relaxation sector decouples in the far past and far future. Using equation (3.42), whose dependence on the relaxation dynamics enters only through $x_{\rm ph}(H)$, we therefore obtain the same asymptotic value of the effective equation of state at both ends of the cosmological evolution,
$\lim_{H\to\infty}w_{\rm eff}=\lim_{H\to 0}w_{\rm eff}=-1+\varepsilon$ (4.18)
In our least-squares fit to the DESI-motivated CPL targets we find $\varepsilon\simeq 0.597$, and hence an asymptotic dark energy equation of state,
$w_{\rm eff}\;\to\;-1+\varepsilon\;\simeq\;-0.403\;\simeq\;-0.40$ (4.19)
@@
both in the early universe and in the infinite future. The bulk viscous contribution in equation (4.3) is therefore a genuinely transient effect that operates only while $x_{\rm ph}(H)$ is of order unity, that is, while the Hubble rate is comparable to the mass gap scale $H_{\star}$. In this intermediate regime the effective equation of state can dip below w=1w=-1 and track the DESI-driven CPL reconstruction, while at very early and very late times the model relaxes back to a regular quintessence-like value $w_{\rm eff}\simeq-0.40$.
4.9 Reproducibility and Ancillary Material
As mentioned earlier, all numbers reported in this section are produced by the ancillary script Statistical Fit.py. The script downloads the public DESI DR1 VAC chain files for $w_{0}w_{a}$CDM for the dataset combination encoded in the base URL, computes the weighted mean and covariance of $(w_{0},w_{a})$, performs the mapping,
$(\varepsilon,\kappa,h_{\star})\;\longrightarrow\;w_{\rm eff}(z)\;\longrightarrow\;(w_{0}^{\rm mod},w_{a}^{\rm mod}),$ (4.20)
and evaluates and minimizes $\chi^{2}$ in equation (4.8). For exact numerical reproducibility, the script writes the file w0wa_compressed.tex and a JSON summary file fit_results.json. The ancillary code file and its outputs are available as a .zip file at (Khan (2026)).
4.10 Phonon Mass Scale associated with $H_{\star}$
In the simplified description of the relaxation time, the characteristic scale $H_{\star}$ can be interpreted as a phonon mass gap via,